Jeff can read 3/4 of a book in 4 hours. How much can he read in 1 hour?
Answer by jim_thompson5910(35256)

3/4 of a book: 4 hours

Divide both parts by 4. This is equivalent to multiplying both parts by 1/4. Let's multiply both sides of the ratio by 1/4

3/4 times 1/4 = (3*1)/(4*4) = 3/16
4 times 1/4 = 4/4 = 1

So the ratio
3/4 of a book: 4 hours
turns into
3/16 of a book: 1 hour


Question: How much can he read in 1 hour?
Answer: 3/16 of a book